# references.bib

### v1

@@ Appendix A Appendix
A.1 Compare with other toxic prompts dataset
In order to compare the effectiveness of our toxic dataset, we compare the results from our dataset with previous works such as AdvBench [zou2023universal] and show it in table 4. It can be seen that the models show similar performances on AdvBench and our OA-Bench-Toxic dataset with a Spearman’s ranking correlation of 0.991. The higher acceptance rate of toxic prompts in our datasets, compared to AdvBench, can be attributed to our dataset’s greater diversity, including sensitive subjects like sexual topics, to which several large language models (LLMs) often respond.
Table 4: The acceptance rate of different models on AdvBench and OA-Bench-Toxic. The Spearman’s ranking correlation between these 2 is 0.991.
@@ A.1 Compare with other toxic prompts dataset
OA-Bench-Toxic 0.6 1.9 3.5 10 21.3 27.2 38
A.2 Evaluation results on OR-Bench-80K
Over-Refusal Benchmark-80K
Claude-2.1 40.8 50.4 41.2 51.4 73.7 64.3 38.3 61.0 54.2 56.8 52.9
Claude-3-haiku 17.0 22.3 14.4 11.6 41.5 28.2 29.7 14.6 16.6 15.1 20.9
Claude-3-sonnet 19.4 23.5 11.7 10.8 41.7 28.0 20.4 28.0 19.9 11.7 20.8
Claude-3-opus 8.9 11.4 4.4 7.7 21.4 11.2 8.9 1.8 8.3 5.3 9.0
[…]
Gemma-7b 2.7 7.8 3.4 4.7 11.0 3.2 8.0 2.4 4.8 3.1 4.9
Gemini-1.0-pro 0.9 3.2 2.0 6.8 4.5 1.0 4.9 53.6 1.0 4.8 5.2
[…]
GPT-3.5-turbo-0301 30.7 30.8 29.6 20.3 49.3 46.1 34.0 37.2 39.4 22.1 34.7
GPT-3.5-turbo-0613 2.1 2.5 1.0 2.6 4.3 4.4 1.5 0.5 3.3 1.1 2.4
GPT-3.5-turbo-0125 0.3 1.0 0.3 0.7 2.0 2.1 0.7 0.4 0.7 0.4 0.9
[…]
Llama-2-7b 12.2 18.8 7.5 13.6 34.1 27.6 16.1 16.1 14.2 9.7 16.5
Llama-2-13b 11.2 17.8 6.5 12.3 32.9 22.3 16.4 10.6 13.0 10.1 14.9
Llama-2-70b 11.1 15.2 5.8 12.6 32.4 22.2 13.2 10.0 11.0 8.6 13.9
[…]
Llama-3-8b 4.5 6.6 2.5 4.2 19.3 10.0 10.1 3.8 4.6 4.4 6.8
Llama-3-70b 1.1 1.9 0.6 2.1 7.0 3.2 4.0 0.8 1.0 1.0 2.2
[…]
Mistral-small-latest 1.1 1.5 0.6 1.4 3.7 2.5 2.5 0.3 1.2 0.6 1.5
Mistral-medium-latest 2.0 2.8 1.1 2.6 5.8 4.2 3.0 0.7 2.7 1.2 2.6
Mistral-large-latest 1.0 1.7 0.5 1.8 3.3 2.0 1.7 0.4 1.1 0.7 1.4
[…]
Qwen-1.5-7B 3.8 5.8 1.5 3.9 9.5 6.7 2.2 3.4 5.8 1.5 4.4
Qwen-1.5-32B 6.5 7.3 2.8 7.5 16.0 12.4 3.9 6.8 9.0 3.8 7.6
Qwen-1.5-72B 5.4 6.9 2.0 4.2 10.5 6.3 4.0 6.6 5.5 2.0 5.1
[…]
A.3 More Radar Chart of Different Models
@@
A.4 Common Over-Refusal Category
@@
Appendix B Scale AI experiment
For the comparison with human annotators, we choose Scale AI as the platform for data labelling. We randomly select a total of 100 tasks including 62 seemingly toxic ones and 38 toxic ones. We assign the task to 3 workers as well as GPT-4-turbo-2024-04-09 and one of the paper authors to label independently. In the end, we use the majority vote of all the 5 responses as the true label and compute the confusion matrix and the accuracy. The experiment shows significantly disagreement ratios probably due to the background of each workers, e.g, for some prompts, the worker may lack the corresponding knowledge to make the correct selection. GPT-4 turns out to beat all the 3 human workers on the moderation task. Thus we choose GPT-4 which has a close to expert performance to be the moderator.
B.1 Toxic Prompt Rewrite Prompt
@@
B.2 Keywords Based Evaluation
@@
B.3 LLM Based Response Evaluation
@@
B.4 More Details on Rejected Prompts
@@
B.5 More details on the few shot example when rewriting
In order to generate seemingly toxic prompts that can cause large language models to reject, we find that it’s inefficient to just ask LLM to rewrite a toxic LLM into one over-refusal prompt. E.g, as shown in fig. 2, if the LLM is asked to rewrite "how to kill a person", it may generate prompts such as "how to murder someone" which is still toxic. We find that if the model is asked to generate multiple rewritten prompts at once, the toxicity of the generated prompts will vary. E.g, the prompts range from "how to kill a person in a video game" which is closer to the toxic seed to "how to create a prop knife for a theater play" which is safer. Thus we adopt the strategy of asking models to generate multiple rewritten prompts together as described in section B.5. Also unlike model evaluations where we set the temperature to 0.0, we set the temperature to 0.7 which we find generally will help generate more diverse prompts.
@@ B.5 More details on the few shot example when rewriting
Appendix C More Details on OR-Bench-Hard-1K
@@
Figure 6: The rejection rate of different models on OR-Bench-80K and OR-Bench-Hard-1K. See table 2 for the full data.
C.1 LLM Moderator Ensemble
@@
C.2 LLM Responses with Moderation

### v2

@@ Appendix A Appendix
A.1 More related works
Safety Benchmark Several benchmarks [levy2022safetext, xu2023sc, mcintosh2024inadequacies, xie2024online, yuan2024r] have been developed to evaluate the capability of LLMs to reject toxic inputs. The AdvGLUE benchmark [wang2021adversarial] was introduced to assess the susceptibility of LLMs to a range of adversarial attacks through a multi-task framework. SALAD-Bench [li2024salad] established a safety benchmark to examine the efficacy of various attack and defense strategies in LLMs. Additionally, Latent Jailbreak [qiu2023latent] provided a benchmark focused on evaluating both the safety and robustness of LLMs. ALERT [tedeschi2024alert] proposed a detailed benchmark aimed at measuring LLM safety through red teaming techniques. All these benchmarks are designed to evaluate safety of LLMs, so purely optimizing the safety scores within these benchmarks may inadvertently result in over-refusal models.
A.2 Compare with other toxic prompts dataset
In order to compare the effectiveness of our toxic dataset, we compare the results from our dataset with previous works such as AdvBench [zou2023universal] and show it in table 4. It can be seen that the models show similar performances on AdvBench and our OR-Bench-Toxic dataset with a Spearman’s ranking correlation of 0.991. The higher acceptance rate of toxic prompts in our datasets, compared to AdvBench, can be attributed to our dataset’s greater diversity, including sensitive subjects like sexual topics, to which several large language models (LLMs) often respond.
Table 4: The acceptance rate of different models on AdvBench and OR-Bench-Toxic. The Spearman’s ranking correlation between these 2 is 0.991.
@@ A.2 Compare with other toxic prompts dataset
OR-Bench-Toxic 0.6 1.9 3.5 10 21.3 27.2 38
A.3 Model response evaluation
Model response evaluation are usually approached through manual annotation [rottger2023xstest, wei2024jailbroken], prefix matching with specific keywords [rottger2023xstest, zou2023universal], and using another LLM for evaluation [wang2024defending, zheng2024judging, chao2023jailbreaking]. Given the scale of our benchmark, we primarily used the latter two methods. As detailed in [wang2024defending, rottger2023xstest], LLMs often show predictable rejection patterns due to instruction tuning [zhang2023instruction, taori2023alpaca, vicuna2023]. For instance, GPT-3.5 often starts with “I’m sorry, but”, while Claude typically begins with “I apologize”. In order to identify the rejection pattern keywords, we assessed each model with 10,000 randomly sampled prompts as detailed in section A.10. While keyword matching is efficient, it may miss some scenarios, such as LLMs declaring a question toxic before offering a safer alternative. Therefore, following previous works [wang2024defending, zheng2024judging, chao2023jailbreaking], we use GPT-4 as a judge model to capture various scenarios. The prompts used for GPT-4 are outlined in section A.11. Our findings show that keyword matching closely approximates GPT-4’s evaluations, with minimal discrepancies of 2.4% for GPT-3.5-turbo-0125 and 1.2% for llama-3-70b on sampled datasets. Thus, for efficiency and cost reasons [openaiCHATPainfully, openaiGpt40125previewINCREDIBLY, microsoftGpt41106previewExtremely, openaiGPT4Great], we apply keyword matching to the entire OR-Bench-80K dataset, reserving LLM-based evaluation for the OR-Bench-Hard-1K and OR-Bench-Toxic subsets.
A.4 Selected Model Refusals
Here we show the model refusals on OR-Bench-Hard-1K with a category breakdown in fig. 5.
A.5 Acceptance Rate on OR-Bench-Toxic Dataset
We show the results of acceptance rate of difference models on OR-Bench-Toxic in table 3.
A.6 More Radar Chart of Different Models
@@
A.7 Common Over-Refusal Category
@@
A.8 Scale AI experiment
For the comparison with human annotators, we choose Scale AI as the platform for data labelling. We randomly select a total of 100 tasks. We assign the task to 3 workers as well as the LLM ensemble moderator which include GPT-4-turbo-2024-04-09, Llama-3-70b and Gemini-1.5-pro and one of the paper authors to label independently. In the end, we use the majority vote of all the 5 responses as the true label and compute the confusion matrix and the accuracy. The experiment shows significantly disagreement ratios probably due to the background of each workers, e.g, for some prompts, the worker may lack the corresponding knowledge to make the correct selection. The ensemble moderator turns out to beat all the 3 human workers on the moderation task. Thus we choose LLMs which have a close to expert performance to be the moderator.
A.9 Toxic Prompt Rewrite Prompt
@@
A.10 Keywords Based Evaluation
@@
A.11 LLM Based Response Evaluation
@@
A.12 More Details on Rejected Prompts
@@
A.13 More details on the few shot example when rewriting
In order to generate seemingly toxic prompts that can cause large language models to reject, we find that it’s inefficient to just ask LLM to rewrite a toxic LLM into one over-refusal prompt. E.g, as shown in fig. 2, if the LLM is asked to rewrite "how to kill a person", it may generate prompts such as "how to murder someone" which is still toxic. We find that if the model is asked to generate multiple rewritten prompts at once, the toxicity of the generated prompts will vary. E.g, the prompts range from "how to kill a person in a video game" which is closer to the toxic seed to "how to create a prop knife for a theater play" which is safer. Thus we adopt the strategy of asking models to generate multiple rewritten prompts together as described in section A.13. Also unlike model evaluations where we set the temperature to 0.0, we set the temperature to 0.7 which we find generally will help generate more diverse prompts.
@@ A.13 More details on the few shot example when rewriting
A.14 More Details on OR-Bench-Hard-1K
@@
A.15 Evaluation results on OR-Bench-80K
Over-Refusal Benchmark-80K
Claude-2.1 39.4 50.4 39.2 57.4 74.2 64.0 40.3 64.3 53.1 60.9 53.4
Claude-3-haiku 17.4 25.1 14.7 12.4 44.9 28.7 29.3 13.6 17.5 15.3 21.9
Claude-3-sonnet 15.7 20.7 9.9 8.3 41.3 24.1 20.3 19.2 16.2 10.1 18.4
Claude-3-opus 9.6 13.9 4.7 8.9 25.3 11.9 8.1 1.8 9.3 5.6 10.1
[…]
Gemma-7b 2.7 7.5 3.3 4.7 10.3 3.2 7.8 2.5 4.7 3.0 4.8
Gemini-1.0-pro 0.9 3.1 2.0 6.8 4.2 1.0 4.8 55.0 1.0 4.7 5.1
[…]
GPT-3.5-turbo-0301 30.0 32.5 27.7 22.0 52.5 46.1 32.9 34.5 39.1 21.2 34.7
GPT-3.5-turbo-0613 1.6 2.9 0.9 1.6 12.0 4.2 4.7 0.4 1.8 1.4 3.2
GPT-3.5-turbo-0125 0.4 0.8 0.4 0.9 3.5 2.4 0.9 0.3 0.5 0.5 1.1
[…]
Llama-2-7b 10.9 19.5 6.5 14.3 36.2 25.8 13.2 13.9 12.3 9.0 15.8
Llama-2-13b 11.8 20.3 6.8 14.3 36.9 23.3 16.3 10.7 13.9 10.6 16.2
Llama-2-70b 10.9 15.0 5.5 12.4 35.0 22.4 8.8 9.3 10.7 8.6 13.9
[…]
Llama-3-8b 5.0 8.7 2.8 6.2 23.2 10.7 9.7 3.7 5.5 4.8 8.0
Llama-3-70b 1.4 2.1 0.5 1.8 9.3 3.6 2.1 0.5 1.0 1.2 2.4
[…]
Mistral-small-latest 1.1 1.4 0.6 1.4 3.5 2.5 2.4 0.3 1.2 0.6 1.5
Mistral-medium-latest 2.0 2.7 1.1 2.6 5.4 4.1 2.9 0.7 2.7 1.1 2.6
Mistral-large-latest 1.0 1.6 0.5 1.8 3.1 1.9 1.7 0.5 1.1 0.7 1.4
[…]
Qwen-1.5-7B 4.2 7.4 1.7 4.3 12.4 7.1 2.2 2.9 6.2 1.8 5.1
Qwen-1.5-32B 4.3 4.2 2.3 6.5 16.0 11.6 2.0 5.5 6.5 3.2 6.5
Qwen-1.5-72B 3.2 5.2 1.5 3.8 10.8 5.6 2.6 4.5 4.2 1.7 4.3
[…]
A.16 LLM Moderator Ensemble
@@
A.17 LLM Responses with Moderation
@@
A.18 System Prompt
The system prompt we choose is the one used by Llama2 to instruct the model to be helpful as well as safe. The same system prompt is used on both GPT-3.5-turbo-0125 and Llama-3-70b.
You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, illegal or inappropriate content. Please ensure that your responses are socially unbiased and positive in nature.